OpenStudy (anonymous):

My 5th Grade son received this: Can some one help please? Thanks Solve for N: 27/36=N/4

OpenStudy (amistre64):

multiply it ll by 4

OpenStudy (anonymous):

so I multiply each by 4?

OpenStudy (anonymous):

Rearrange so N=(27/36)*4

OpenStudy (anonymous):

is this proper for a 5th grade kid?

OpenStudy (amistre64):

\[\frac{N}{4}*4=N*\frac{4}{4}=N*1=N\]

OpenStudy (amistre64):

5th grade is fine for teaching how to use math operations,

OpenStudy (anonymous):

It is basic algebra.

OpenStudy (anonymous):

ok thanks